# Meritocracy and Its Discontents: Long-run Effects of Repeated School Admission Reforms

### v1

@@ 4 Long-run Impacts
To assess long-run effects of the meritocratic centralization, we provide two sets of empirical analysis. First, we investigate the persistence of the distributional effect of centralization that we document in the short-run analysis. We do so by a difference-in-differences strategy comparing labor-market outcomes of urban- and rural-born individuals across birth cohorts with differential exposure to the centralized admissions. Our analysis shows that the centralization increased both (a) the number of career elites born in urban areas and (b) the number of career elites living in urban areas as adults. The admission reforms thus influence the regional origins and destinations of career elites.
Second, we explore the national productive efficiency of the two admission systems. We find that cohorts more intensely exposed to the centralized admissions produced a larger national number of civil officials who were promoted to top ranks. This finding indicates that the centralization boosted the national production of career elites, at least for the specific important category of top-ranking government officials.
4.1 Regional Distribution of Career Elites
@@
To analyze long-run effects of the centralization on students’ career outcomes, we use the Personnel Inquiry Records (PIR) published in 1939.252525Japan went into the military regime with the promulgation of the National General Mobility Act in 1938. Although the military government began massive social interventions, their effect was relatively small in 1939 (Moriguchi and Saez (2008)). In addition, most of the information in PIR (1939) is based on 1938 data. The PIR is an equivalent of Who’s Who, which compiles a selective list of socially distinguished individuals (including high-income earners, top business managers, elite professionals, and high-ranking public servants) and provides their biographical information. In total, the 1939 PIR lists 55,742 individuals or 0.15% of the adult Japanese population of that time. We detail the selection process in Appendix Section A.4.1.
@@
To capture the effects of the first period of the centralized admission system in 1902–1907, we use the cohorts born in 1880–1894, who turned 17 years old (the age eligible for application) in 1897–1911. The cohorts born in 1880–1894 were 45 to 59 years old in 1939. The number of individuals listed in the PIR in each of these cohorts is about 1,800. We use the following information from the PIR data for each individual: full name, birth year, birth prefecture, prefecture of residence, final education, occupation titles and organization names, the medal for merit and the court rank awarded (if any), and the amount of national income tax and business tax paid (if any).
@@
We define the following groups of elites as a subset of PIR-listed individuals: (1) the top 0.01% and 0.05% income earners according to the national income distribution, (2) prestigious medal recipients (civilians who received either the medal of the fifth order of merit or above, or the court rank of the junior fifth rank or above), (3) corporate executives (individuals who hold an executive position in a corporation and pay a positive amount of income or business tax), (4) top politicians and bureaucrats (individuals whose occupation is either Imperial Diet member, Cabinet member, or high-ranking central government official), and (5) Imperial University professors or associate professors. Appendix Section A.4.1 and Table A.8 provide detailed definitions and descriptions of each group. These categories encompass economic, social, cultural, political, and academic definitions of career elites.
@@
We use this data to count the number of elites in each group defined by birth prefecture and birth cohort. These counts allow us to conduct a difference-in-differences analysis that compares long-term career outcomes of urban- and rural-born individuals by each cohort’s exposure to the centralized admission system. Descriptive statistics of main variables are summarized in Appendix Table A.2.
@@
Urban-Rural Disparity in Producing Career Elites
We estimate the long-run impacts of the centralized admission system by conducting a difference-in-differences analysis by birth cohorts and birth areas. The key idea behind our empirical strategy is that applicants born in the Tokyo area should experience a greater gain in entering Schools 1–8 under Capp relative to Dapp, since the centralized system is designed to be more meritocratic and high-achieving students are disproportionately located in urban areas. This expectation is confirmed in Figure 2 and Table 1. We exploit this differential gain in school access to compare the career outcomes of individuals born inside and outside the Tokyo area by the cohort’s exposure to Capp. If admission to Schools 1–8 increases one’s chance of becoming a career elite, we should observe a greater number of elites born inside the Tokyo area for the cohorts exposed to Capp. We estimate a difference-in-differences specification as follows:
@@
where $Y_{pt}$ is the number of elites born in cohort $t$ and prefecture $p$. $Centralized_{t}$ is a measure of cohort $t$’s exposure to Capp, which is, in the baseline specification, a binary indicator that cohort $t$ turned 17 during Capp (1902–1907). $Tokyo\_area_{p}$ is the indicator that takes 1 if prefecture $p$ is in the Tokyo area. $\gamma_{p}$ and $\gamma_{t}$ are prefecture and cohort fixed effects. To allow for serial correlation of $\epsilon_{pt}$ within prefecture over time, we cluster the standard errors at the prefecture level in our baseline specification. [^26] Bertrand, Duflo and Mullainathan (2004) evaluate approaches to deal with serial correlation within each cross-sectional unit in panel data. They suggest that clustering the standard errors on each cross-section unit performs well in settings with 50 or more cross-section units, as in our setting. In addition, we report the results of clustering at cohort level, which are estimated by wild cluster bootstrap (Cameron and Miller, 2015; Roodman et al., 2019) due to the small number of clusters (15 cohorts).
@@
The above regression defines $Centralized_{t}$ to be a binary indicator, as the simplest proxy for the intensity of exposure to Capp. In reality, however, a nontrivial number of unsuccessful applicants retook the exam at age 18 and beyond. As a result, the cohorts who turned age 17 in 1899–1901 were partially and increasingly exposed to Capp (as some of them took the exam in 1902). The cohorts who turned age 17 in 1902–1904 were fully exposed to Capp. The cohorts who turned age 17 in 1905–1907 were partially and decreasingly exposed to Capp (as some of them took the exam in 1908). Finally, the intensity of exposure drops to zero for the cohorts who turned age 17 in 1908.272727The intensity of exposure to Capp is estimated to be 0.00, 0.02, 0.09, 0.36, 1.00, 0.98, 0.91, 0.64, and 0.00 for the cohort who turned age 17 in 1896–98, 1899, 1900, 1901, 1902–04, 1905, 1906, 1907, and 1908–11, respectively. See Appendix Section A.4.1 for data and methods. In the following visual analysis, we explicitly incorporate cohort’s intensity of exposure to Capp by coloring cohorts according to their intensity of exposure to Capp.
@@
We first check whether the number of Imperial University graduates born inside the Tokyo area increased for the cohorts exposed to Capp. Since all Schools 1–8 graduates were automatically admitted to an Imperial University during this period, the areas that produced more Schools 1–8 entrants should produce more Imperial University graduates. Figure 3 confirms this expectation. Panel (a) compares the average number of Imperial University graduates who were born in prefectures inside and outside the Tokyo area by cohorts (represented by their birth year plus 17 on the horizontal axis). Figure 3 (b) shows that the urban-rural difference in the number of Imperial University graduates rises as the intensity of exposure to Capp increases and falls sharply when Capp ends in 1908. Column (1) in Table 2 shows that the estimate of $\beta$ in the above regression is positive and statistically significant.
@@
Our main results are presented in Figure 3 (c)–(f) and Table 2 Columns (2)–(7). Figure 3 (c)–(f) show difference-in-differences plots that compare the number of elites (the top 0.05% income earners and medal recipients) who were born inside and outside the Tokyo area by the cohort’s exposure to Capp. For both elite categories, the plots show that the difference between the Tokyo area and the rest grows as the intensity of exposure to Capp increases, and then drops sharply when Capp ends. Appendix Figure A.7 shows qualitatively similar difference-in-differences plots for the top 0.01% income earners, corporate executives, top politicians and bureaucrats, and Imperial University professors.
@@
Table 2 Columns (2)–(7) show that the long-run effects of Capp are economically and statistically significant. For the cohorts exposed to Capp, the number of career elites born inside the Tokyo area (compared to those born outside the Tokyo area) increases by 34% for the top 0.01% income earners, 24% for the top 0.05% income earners, 36% for medal recipients, 14% for corporate executives, 50% for top politicians and bureaucrats, and 47% for Imperial University professors (in Panel B). [^28] In Appendix Table A.10, we also examine other groups of elite professionals and obtain similar results for scholars, engineers, and physicians. In Panel A, we control only for cohort and prefecture fixed effects. In Panel B, we additionally control for time- and cohort-varying prefecture characteristics. [^29] Cohort birth population, the number of primary schools, the number of middle school graduates, and prefecture-level GDP. See Appendix Section A.4.1 for variable definitions and data sources. The coefficients fall slightly in magnitude after adding control variables, but remain sizable.
@@
Table 2 Panel C shows that the effects are symmetric with respect to the direction of the admission reforms, i.e., the change from Dapp to Capp and the change from Capp to Dapp produce similar effects of the opposite sign. In Table 2 Panel D, we replace the centralization dummy by the cohort’s intensity of exposure to Capp. The results remain qualitatively the same as the baseline results.
@@
These results suggest that almost four decades after its implementation, Capp had lasting effects on the career trajectories of students. The above results are robust to alternative specifications. First, the analysis in Table 2 Panel D assumes that the cohort’s intensity to exposure to Capp is exogenous and the same across years, which may be a strong assumption. However, even when we drop the cohorts who are exposed to both Capp and Dapp (i.e., cohorts who became age 17 in 1901 and 1907) from the sample, we still find qualitatively the same results (Appendix Table A.11). Second, we test if the assumption of parallel pre-event trends holds. Appendix Table A.12 verifies that the differences in pre-event trends between the areas of comparison are small and mostly insignificant.
@@
Another potential threat to our identification strategy is that there may be some age-specific trends in the number of elites that covary with the cohort-region variation we use. Specifically, the number of elites selected in the 1939 long-term data peaks at around the cohort who were 51 years old in 1939 (corresponding to the cohort who turned age 17 in 1905) and gradually falls for younger and older cohorts, suggesting that there are certain ages at which individuals are more likely to be listed in the long-term data. Such age effects may generate different trends in the number of elites born in the Tokyo and other areas, due, for example, to differences in population size across these areas. To address this concern, we use an earlier edition of the PIR published in 1934, construct the prefecture-cohort level data for the same cohorts used in our main analysis (but observed 5 years earlier), and conduct similar regression analyses. The results in Appendix Table A.13 confirm that our key results remain qualitatively the same even when we use the 1934 PIR data.
@@
Finally, we conduct placebo tests to examine if the results are driven by other factors such as the sample selection of the PIR or changes in cohort populations. The urban-rural difference in the cohort’s birth populations does not change significantly with the cohort’s exposure to Capp, as shown in Appendix Table A.14. As an additional placebo test, we also look at unrelated career outcomes. Among the distinguished individuals listed in the PIR, we expect that landlords (defined in Appendix Section A.4.1) are least likely to be affected by the introduction of Capp as receiving higher education was not a typical pathway to become a landlord. As shown in Appendix Table A.14, the estimated effect of Capp on the number of landlords is small and statistically insignificant. These findings further support our conclusion that the meritocratic centralization enlarged the urban-rural gap in the capacity to produce future career elites.
Geographical Destinations of Career Elites
@@
Having established that Capp affects the geographic origins of elites, we now ask how it affects their geographic destinations. While the former is about regional inequality in educational opportunities, the latter is about regional inequality in the supply of highly skilled human capital, which potentially affects economic growth and inequality.
@@
We first examine if the centralization-induced increase in inter-regional mobility in the short run boosted the geographical mobility of elites in the long run. Surprisingly, it did not: The urban-rural difference in the fraction of elites whose prefectures of residence differ from their birth prefectures did not increase under Capp, as shown in Appendix Table A.14. We find similar results when we use the distance between an elite’s birth prefecture and his prefecture of residence as an alternative measure of long-run mobility. This result suggests that, even though a greater number of students born in the Tokyo area entered rural schools under Capp, most of them returned to the Tokyo area when pursuing their careers
@@
If the greater number of Tokyo-area born students admitted to rural schools under Capp returned to the Tokyo area eventually for their subsequent careers, we should observe a greater number of elites living in the Tokyo area for the cohorts exposed to Capp. To test this hypothesis, we redefine the outcome variables by changing the prefecture ($p$) from birth prefecture to prefecture of residence and estimate the same equation.
@@
Table 3 shows large positive effects of Capp on the urban-rural gap in the number of elite residents. For the cohorts exposed to Capp relative to Dapp, the number of elites living in the Tokyo area in their middle age (compared to those living outside the Tokyo area) increases by 28% for the top 0.01% income earners, 23% for the top 0.05% income earners, 29% for medal recipients, 22% for corporate executives, 20% for top politicians and bureaucrats, and 50% for Imperial University professors (Panel B). These results suggest that the meritocratic centralization intensified the concentration of career elites in urban areas in the long run.
4.2 National Production of Career Elites

### v2

@@ 4 Long-run Impacts
To assess long-run effects of meritocratic centralization, we provide two sets of empirical analysis. First, we investigate the persistence of the distributional effect of centralization that we document in the short-run analysis. We do so by a difference-in-differences strategy comparing labor-market outcomes of urban- and rural-born individuals across birth cohorts with differential exposure to the centralized admissions. Our analysis shows that centralization increased both (a) the number of occupational elites born in urban areas and (b) the number of occupational elites living in urban areas as adults. The admission reforms thus influence the regional origins and destinations of elites.
Second, we compare the national production of elites between the two admission systems. We find that cohorts more intensely exposed to the centralized admissions produced a larger national number of civil officials who were promoted to top ranks, as well as other types of occupational elites (such as top income earners). This finding indicates that centralization boosted the national production of individuals reaching leadership positions.
4.1 Regional Distribution of Elites
@@
To analyze the long-run effects of centralization on students’ career outcomes, we use the Personnel Inquiry Records (PIR) published in 1939.262626Japan went into the military regime after the National General Mobility Act of 1938. Although the military government began massive economic interventions, their effect was relatively small in 1939 (Moriguchi and Saez, 2008). Moreover, most of the information in the 1939 PIR was based on 1938 data. The PIR is an equivalent of Who’s Who, which compiles a highly selective list of socially distinguished individuals—encompassing economic, political, cultural, and aristocratic elites—and provides their biographical information. The 1939 PIR lists 55,742 individuals or 0.15% of the adult Japanese population of that time. [^27] We discuss the PIR’s selection criteria in Appendix Section A.4.1. According to their occupational distribution in Appendix Table A.8 Column (1), 66% hold managerial positions, 7.5% are politicians or bureaucrats, and the rest consists mainly of professionals such as scholars and engineers. Throughout this section, we refer to the individuals listed in the PIR as “elites.”
@@
To capture the effects of the first period of the centralized admissions in 1902–1907, we use the cohorts born in 1880–1894, who turned 17 years old (the age eligible for application) in 1897–1911. These cohorts were 45 to 59 years old in 1939. The number of elites in each of these cohorts is about 1,800. We use the following information from the PIR for each individual: full name, birth year, birth prefecture, prefecture of residence, final education, occupation titles and organization names, the medal for merit and the court rank awarded (if any), and the amount of national income tax and business tax paid (if any). Among all elites in our sample, 23% graduated from Imperial Universities.
@@
We define the following subgroups of elites (as a subset of PIR-listed individuals): (1) the top 0.01% and 0.05% income earners according to the national income distribution, (2) prestigious medal recipients (civilians who received high-ranking imperial medals for their exceptional service or merit in various fields), [^28] These prestigious medal recipients held a broad range of occupations, including scholars, managers, engineers, and bureaucrats as shown in Appendix Table A.8 Column (4). (3) corporate executives (individuals who hold an executive position in a corporation and pay a positive amount of income or business tax), (4) top politicians and bureaucrats (individuals whose occupation is either Imperial Diet member, Cabinet member, or high-ranking central government official), (5) Imperial University professors or associate professors, (6) all occupational elites (all individuals listed in the PIR except for hereditary elites defined as individuals whose sole occupation is either peerage or landlord). Appendix Section A.4.1 and Table A.9 provide detailed definitions and descriptions of each group. These categories encompass economic, political, and cultural definitions of occupational elites.
@@
We use this data to count the number of elites in each subgroup defined by birth prefecture and birth cohort. These counts allow us to conduct a difference-in-differences analysis that compares long-term career outcomes of urban- and rural-born individuals by each cohort’s exposure to the centralized admission system. Descriptive statistics of main variables are summarized in Appendix Table A.2.
@@
Urban-Rural Disparity in Producing Elites
We estimate the long-run impacts of the centralized admission system by conducting a difference-in-differences analysis by birth cohorts and birth areas. The key idea behind our empirical strategy is that urban-born applicants experienced a greater gain in entering Schools 1–8 under centralization relative to decentralization, as shown in the short-run analysis in Figure 2 and Table 1. We exploit this differential gain in school access to compare the career outcomes of individuals born inside and outside the Tokyo area by the cohort’s exposure to centralization. If admission to Schools 1–8 increases one’s chance of becoming an elite, we should observe more elites born inside the Tokyo area for the cohorts exposed to centralization. We estimate a difference-in-differences specification as follows:
@@
where $Y_{pt}$ is the number of elites born in cohort $t$ and prefecture $p$. $Centralized_{t}$ is a measure of cohort $t$’s exposure to centralization, which is, in the baseline specification, a binary indicator that cohort $t$ turned 17 during centralization (1902–1907). $Tokyo\_area_{p}$ is the indicator that takes 1 if prefecture $p$ is in the Tokyo area. $\gamma_{p}$ and $\gamma_{t}$ are prefecture and cohort fixed effects. To allow for serial correlation of $\epsilon_{pt}$ within prefecture over time, we cluster the standard errors at the prefecture level in our baseline specification. [^29] Bertrand, Duflo and Mullainathan (2004) evaluate approaches to deal with serial correlation within each cross-sectional unit in panel data. They suggest that clustering the standard errors on each cross-section unit performs well in settings with 50 or more cross-section units, as in our setting. In addition, we report the results of clustering at cohort level, which are estimated by wild cluster bootstrap (Cameron and Miller, 2015; Roodman et al., 2019) due to the small number of clusters (15 cohorts).
@@
The above regression defines $Centralized_{t}$ to be a binary indicator as the simplest proxy for the intensity of exposure to centralization. In reality, however, a nontrivial number of unsuccessful applicants retook the exam at age 18 and beyond. As a result, the cohorts who turned age 17 in 1899–1901 were partially and increasingly exposed to centralization (as some of them took the exam in 1902). The cohorts who turned age 17 in 1902–1904 were fully exposed to centralization. The cohorts who turned age 17 in 1905–1907 were partially and decreasingly exposed to centralization (as some of them took the exam in 1908). Finally, the intensity of exposure drops to zero for the cohorts who turned age 17 in 1908.303030The intensity of exposure to centralization is estimated to be 0.00, 0.02, 0.09, 0.36, 1.00, 0.98, 0.91, 0.64, and 0.00 for the cohort who turned age 17 in 1896–98, 1899, 1900, 1901, 1902–04, 1905, 1906, 1907, and 1908–11, respectively. See Appendix Section A.4.1 for data and methods. In the following visual analysis, we explicitly incorporate cohort’s intensity of exposure to centralization by coloring cohorts according to their intensity of exposure to centralization.
@@
We first check whether the number of Imperial University graduates born inside the Tokyo area increased for the cohorts exposed to centralization. Since all Schools 1–8 graduates were automatically admitted to an Imperial University during this period, the areas that produced more Schools 1–8 entrants should produce more Imperial University graduates. Figure 3 confirms this expectation. Panel (a1) compares the average number of Imperial University graduates who were born inside and outside the Tokyo area by cohorts (represented by their birth year plus 17 on the horizontal axis). Panel (a2) shows that the urban-rural difference in the number of Imperial University graduates rises as the intensity of exposure to centralization increases and falls sharply when centralization ends in 1908. Column (1) in Table 2 shows that the estimate of $\beta$ in the above regression is positive and statistically significant.
@@
Our main results are presented in Figure 3 Panels (b)–(d) and Table 2 Columns (2)–(8). Figure 3 Panels (b1)–(d1) show difference-in-differences plots that compare the number of elites (the top 0.05% income earners, medal recipients, and all occupational elites) who were born inside and outside the Tokyo area by the cohort’s exposure to centralization. For these elite categories, Panels (b2)–(d2) show that the difference between the Tokyo area and the rest grows as the intensity of exposure to centralization increased, and then drops sharply when centralization ended. Appendix Figure A.7 shows qualitatively similar difference-in-differences plots for the top 0.01% income earners, corporate executives, top politicians and bureaucrats, and Imperial University professors.
@@
Table 2 Columns (2)–(7) show that the long-run effects of centralization are economically and statistically significant. For the cohorts exposed to centralization, the number of elites born inside the Tokyo area (compared to those born outside the Tokyo area) increases by 34% for the top 0.01% income earners, 24% for the top 0.05% income earners, 36% for medal recipients, 13% for corporate executives, 50% for top politicians and bureaucrats, 38% for Imperial University professors, and 14% for all occupational elites (in Panel B). [^31] In Appendix Table A.11, we also examine other types of occupational elites and obtain similar results for scholars, engineers, and physicians. In Panel A, we control only for cohort and prefecture fixed effects. In Panel B, we additionally control for time- and cohort-varying prefecture characteristics. [^32] Cohort birth population, the number of primary schools, the number of middle school graduates, and prefecture-level GDP. See Appendix Section A.4.1 for variable definitions and data sources. The coefficients fall slightly in magnitude after adding control variables, but remain sizable.
@@
Table 2 Panel C shows that the effects are symmetric with respect to the direction of the admission reforms, i.e., the change from decentralization to centralization and the change from centralization to decentralization produce similar effects of the opposite sign. In Table 2 Panel D, we replace the centralization dummy by the cohort’s intensity of exposure to centralization. The results remain qualitatively the same as the baseline results.
@@
These results suggest that almost four decades after its implementation, centralization had lasting effects on the career trajectories of students. The above results are robust to alternative specifications. First, the analysis in Table 2 Panel D assumes that the cohort’s intensity to exposure to centralization is exogenous and the same across years, which may be a strong assumption. However, even when we drop the cohorts who are exposed to both centralization and decentralization (i.e., cohorts who became age 17 in 1901 and 1907) from the sample, we still find qualitatively the same results (Appendix Table A.12). Second, we test if the assumption of parallel pre-event trends holds. Appendix Table A.13 verifies that the differences in pre-event trends between the areas of comparison are small and mostly insignificant.
@@
Another potential threat to our identification strategy is that there may be some age-specific trends in the number of elites that covary with the cohort-region variation we use. Specifically, the number of elites listed in the 1939 PIR data peaks at around the cohort who were 51 years old in 1939 (corresponding to the cohort who turned age 17 in 1905) and gradually falls for younger and older cohorts. This suggests that there are certain ages at which individuals are more likely to be listed in the long-term data. Such age effects may generate different trends in the number of elites born in the Tokyo and other areas, due, for example, to differences in population size across these areas. To address this concern, we use an earlier edition of the PIR published in 1934, construct the prefecture-cohort level data for the same cohorts used in our main analysis (but observed 5 years earlier), and conduct similar regression analyses. The results in Appendix Table A.14 confirm that our key results remain qualitatively the same even when we use 1934 PIR data.
@@
Finally, we conduct placebo tests to examine if the results are driven by other factors such as the sample selection of the PIR or changes in cohort populations. The urban-rural difference in the cohort’s birth populations does not change significantly with the cohort’s exposure to centralization, as shown in Appendix Table A.15. As an additional placebo test, we also look at unrelated career outcomes. Among the distinguished individuals listed in the PIR, we expect that landlords (defined in Appendix Section A.4.1) are least likely to be affected by the introduction of centralization as receiving higher education was not a typical pathway to become a landlord. As shown in Appendix Table A.15, the estimated effect of centralization on the number of landlords is small and statistically insignificant. These findings further support our conclusion that meritocratic centralization enlarged the urban-rural gap in the capacity to produce future individuals in leadership positions.
Geographical Destinations of Elites
@@
Having established that centralization affects the geographic origins of occupational elites, we now ask how it affects their geographic destinations. While the former is about regional inequality in educational opportunities, the latter is about regional inequality in the supply of highly skilled human capital.
@@
We first examine if the centralization-induced increase in inter-regional mobility in the short run boosted the geographical mobility of elites in the long run. It did not: The urban-rural difference in the fraction of elites whose prefectures of residence differ from their birth prefectures did not increase under centralization, as shown in Appendix Table A.15. We find similar results when we use the distance between an elite’s birth prefecture and his prefecture of residence as an alternative measure of long-run mobility. This result suggests that, even though a greater number of students born in the Tokyo area entered rural schools under centralization, most of them returned to the Tokyo area when pursuing their careers.
@@
If the increased number of Tokyo-area-born students admitted to rural schools under centralization returned to the Tokyo area eventually for their careers, we should observe a greater number of elites living in the Tokyo area for the cohorts exposed to centralization. To test this hypothesis, we redefine the outcome variables by changing the prefecture ($p$) from birth prefecture to prefecture of residence and estimate the same equation.
@@
Table 3 shows large positive effects of centralization on the urban-rural gap in the number of elite residents. For the cohorts exposed to centralization relative to decentralization, the number of elites living in the Tokyo area in their middle age (compared to those living outside the Tokyo area) increases by 30% for the top 0.01% income earners, 25% for the top 0.05% income earners, 30% for medal recipients, 23% for corporate executives, 19% for top politicians and bureaucrats, 41% for Imperial University professors, and 20% for all occupational elites (Panel B). These results suggest that meritocratic centralization intensified the concentration of elites in urban areas in the long run.
4.2 National Production of Elites